OpenStudy (swissgirl):

Show that any non-empty finite set S ⊂ ℝ contains both its supremum and infimum. (Hint: use induction)

OpenStudy (swissgirl):

@abb0t

OpenStudy (swissgirl):

To be honest I really didnt study the chapter fully yet I was just looking for a tutor so hopefully I wont be lost lol

OpenStudy (abb0t):

\[S \subseteq R\] such that |S| = n Proceed by induction on n. If n = 1, then S = {a} for some \[a \in R\]. Then, sup S = a. Assume that all subsets of R of order k contain their respective suprema. Let S be a subset of R in order of k+1. Label the elements of S as: \[\left\{ a_1, a_2...a_{k+1} \right\}\] Let \[T = \left\{ a_1, a_2...a_k \right\} \] By the inductive assumption: \[\sup T= a_{i_o}\] for some \[1 \le i_o \le k\] and \[\sup(S-T) = a_{k+1}\] Thus clamining case 1: \[a_{i_o} \le a_{k+1}\] since \[a_{i_o} = \sup T\] and \[a_{i_o} \le a_{k+1}\], you know that for all \[1 \le i_o \le k, a_i \le a_{i_o} \le a_{k+1}\] thus \[\sup S = a_{k+1}\] Case 2: \[a_{i_o} \ge a_{k+1}\] since \[a_{i_o} = \sup T\] and \[a_{i_o} \ge a_{k+1} \] you know that for all \[1 \le i \le k+1, a_i \le a_{i_o}\] thus \[\sup S = a_{i_o}\] Thus, S contains its supremum

OpenStudy (swissgirl):

Give me time to read it. Thanks :)

OpenStudy (swissgirl):

Wow this is great :)

OpenStudy (swissgirl):

I didnt follow case 2 though

OpenStudy (swissgirl):

Ummm Also werent we supposed to prove that it contains its supremum and infimum

OpenStudy (swissgirl):

Like I followed what u did but i am just confused why we needed to show case 2

OpenStudy (abb0t):

Well supremem is the upper bound, and infimum is the lower bound. I just used the definitions.

OpenStudy (swissgirl):

No like how did u show S contains the infimum

OpenStudy (abb0t):

Case 2.

OpenStudy (abb0t):

I showed4 supremum.

OpenStudy (swissgirl):

Ohhhhhh ok I thought i was going senile

OpenStudy (abb0t):

:) nah don't say that.

OpenStudy (swissgirl):

lol Thankssssssss. THIS WAS AWESOME. U R REALLY CLEARRRRRRRR